OpenStudy (anonymous):

simplify negative square root of 189

OpenStudy (anonymous):

complex numbers wud help u

OpenStudy (anonymous):

First, I would factor 189. See if you can find a perfect square factor.

OpenStudy (anonymous):

do u know what is square root of -1

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

@nancydiaz14 Have you learned about imaginary numbers yet?

OpenStudy (anonymous):

there is a concept of i for root -1 hv u heard of it

OpenStudy (anonymous):

no we barely started learning about square roots.

OpenStudy (anonymous):

I think I am supposed to find the factor of 189

OpenStudy (anonymous):

I would think that simplify means to factor out any perfect squares. For example: \(\sqrt{12} = 2\sqrt{3}\) because 12 factors to 4 and 3. \(\sqrt{4} = 2\).

OpenStudy (anonymous):

so could I factor it with 9 * 21 and then get -3 times the square root of 21?

OpenStudy (anonymous):

That's where I'm not sure of. Not having been in your class, I can't be sure what your teacher is asking for. \(\sqrt{189} = \pm 3\sqrt{21}\). I think you are correct with the -3.